An RL circuit has an emf of 20 volts, a resistance of 20 ohms, an inductance of 5 H with an initial current of 4 mA. What is the current after 0.05 seconds?



1
Expert's answer
2021-12-17T14:04:27-0500

Differential equation of a LR circuit 


"Li'+Ri=U"

Given "U=20V, R=20\ Ohm, L=5\ H"

Substitute


"5i'+20i=20""i'+4i=4"

Integrating factor


"\mu(t)=e^{\int4dt}=e^{4t}"

"e^{4t}(i'+4i)=4e^{4t}"

"d(e^{4t}i)=4e^{4t}dt"

Integrate


"\int d(e^{4t}i) =\int 4e^{4t}dt"

"e^{4t}i=e^{4t}+C"

"i(t)=1+Ce^{-4t}"

"i(0)=4mA=4\times 10^{-3} A"

"4\times 10^{-3} =1+C"

"C=-0.996"

"i(t)=1-0.996e^{-4t}"

"i(0.05)=1-0.996e^{-4(0.05)}"

"i(0.05)=184.5\ mA"